OpenStudy (anonymous):

Find the center and radius of the circle whose equation is given. x^2+y^2+8x-9=0

OpenStudy (anonymous):

Complete the square on the x part!

OpenStudy (anonymous):

what is the final answer

OpenStudy (anonymous):

That gives you \[ (x+4)^2 - 16 + y^2 - 9 = 0 \]. Simplified it is \[(x+4)^2+y^2=25=5^2\]

OpenStudy (anonymous):

The center is at (-4,0) and the radius is 5.